OpenStudy (lamechildidc):

Okay same thing, I have an idea of the answer But i am not sure. Given the exponential equation 3^x = 27, what is the logarithmic form of the equation in base 10? x = log base 10 of 3, all over log base 10 of 27 x = log base 10 of 27, all over log base 10 of 3 x = log base 2 of 3, all over log base 2 of 27 x = log base 2 of 10, all over log base 2 of 3

OpenStudy (lamechildidc):

@galaxystare

OpenStudy (lamechildidc):

@Roxy.girl

OpenStudy (lamechildidc):

I put B

OpenStudy (anonymous):

x = log base 10 of 3, all over log base 10 of 27 x=log103log1027 x = log base 10 of 27, all over log base 10 of 3 log1027log103 x = log base 2 of 3, all over log base 2 of 27 x = log base 2 of 10, all over log base 2 of 3

OpenStudy (anonymous):

so i would say b ya

OpenStudy (lamechildidc):

r u sure haha this is my seg exam i have to pass

OpenStudy (lamechildidc):

@MrNood what do you think

OpenStudy (roxy.girl):

Yup your right :)

OpenStudy (lamechildidc):

bless :') thank you how about this? What are the explicit equation and domain for a geometric sequence with a first term of 4 and a second term of −8? an = 4(−2)n − 1; all integers where n ≥ 0 an = 4(−2)n − 1; all integers where n ≥ 1 an = 4(−12)n − 1; all integers where n ≥ 1 an = 4(−12)n − 1; all integers where n ≥ 0

OpenStudy (lamechildidc):

i think its B

OpenStudy (mrnood):

\[\log a ^{m} = m \log a\] Just start with 3^x use my equation to write log 3^x in another way

OpenStudy (mathmale):

Without explanations of WHY you think this or that, "I think it's B" responses make very little sense.

OpenStudy (lamechildidc):

Because base of 10 and 3^x=27 is also log10 27=x

OpenStudy (mathmale):

3^x = 27 is in exponential form. Take the common log of both sides of this equation, applying rules of logs. Show your results, please.

OpenStudy (mathmale):

log (3^x) = log 27. Solve for x, please.

OpenStudy (lamechildidc):

i dont understand

OpenStudy (lamechildidc):

@mathmale x = 3?

OpenStudy (mathmale):

Please separate what you understand and what you don't. Then I'd be happy to provide further explanations. Yes, x=3 is correct, but I'd like to know how you arrived at that.

OpenStudy (lamechildidc):

because it 3 x 3 x 3 = 27

OpenStudy (mathmale):

that's correct, but it does not follow the instructions that came with this problem: "what is the logarithmic form of the equation in base 10?" You must apply common logs to arrive at the correct answer. Please note that I have already given you hints regarding what to do; review this conversation, please.

OpenStudy (lamechildidc):

Wait. I already have the answer to this. Can you help with another if i post a different question?